EFI operational status:
The station operator visited the station and replaced the Episensor. All channels from the Episensor strong motion sensor were non-seismic day 2018,061-02:03 to 270-19:49.
IDA engineers visited the station and replaced the KS54000 seismometer with a Trillium T360 seismometer, installed a Hyperion infrasound sensor, new station computers, and a weather transmitter.
An IDA engineer visited the station and replaced the CMG-3T seismometer with a Trillium120PH seismometer. Installation of a Setra microbarograph and power equipment were also completed.
The masses for the CMG-3T seismometer were saturated following a short outage at the station. All channels from the CMG-3T were non-seismic/saturated from day 2012,307-01:07 to 2012,311-00:20 until the sensor could be unlocked then re-centered.
An IDA engineer visited the station and upgraded the solar power system.
The station is back on-line as of day 2011,156-02:21 after an IDA engineer cleared the ISP and restarted data acquisition.
The station is down since about day 2011,144-12:38 due to excessive gapping of data. Currently telemetry to the station is also down. The station operator has been requested to visit the station to investigate.
The station operator replaced the efi00 Q330 digitizer. All BB*00 channels from the KS-54000 seismometer were restored at 2010,060-15:10.
The station is back on-line as of day 2010,035-16:42 after the station operator cycled power to the CMG-3T seismometer, and an IDA engineer centerered the masses for the CMG-3T seismometer. All channels from the KS-54000 seismometer remain non-seismic due to the faulty Q330 efi00 digitizer.
IDA station EFI is down as of day 2010,012-13:57 due to problems with the Q330 efi00 digitizer that are yet to be determined.
IDA station EFI has been upgraded to a Quanterra Q330HR system. Seismometers include a KS-54000 and CMG-3T, and an episensor strong motion detector.
the ARS, the GPS clock, and cycled power to the DAS.
The station is down as of about day 2008,163 due to problems with the ARS that are yet to be determined.
The BH1-00,LH1-00,and VH1-00 channels from the KS-54000 seismometer were restored as of about day 2008,087 after the station operator replaced a faulty circuit board in the DAS.
The BH1-00,LH1-00,and VH1-00 channels from the KS-54000 seismometer are not being acquired since about day 2008,063-17:47 for reasons yet to be determined.
The station was down from about day 2007,268-02:23 to 2007,275-19:20 due to communication problems between the ISP and the ARS.
The station is back online as of about day 072-13:33 after the station operator cycled power to the ARS.
The station is down as of about day 069-01:12 due to communication problems between the DAS and the ISP.
IDA station EFI was down from day 042-10:50 to 045-13:26 due to communication problems betweeen the DAS and the ISP.
IDA station EFI was down from about day 024 to 030 due to communication problems between the ISP and the ARS.
An IDA engineer visited the station to upgrade the telemetry system, and replace the Main Power Supply and the FBA-23 accelerometer.
The East-West channels for the CMG-3T seismometer were restored at approximately day 256-18:50 (Sept.13) after the station operator recentered the masses for the CMG-3T seismometer.
The East-West channels for the CMG-3T seismometer are out as of day 219 for reasons yet to be determined.
The vertical channels are back as of day 215-14:00 after the station operator rebooted the DAS.
The vertical channels for the KS-54000 seismometer are out as of day 209-15:09 for reasons yet to be determined.
The station operator installed the new tape drive and data recording has resumed.
The tape drives at the station have failed. A new tape drive is being tested and will be sent to the station.
An IDA engineer visited the site to survey and check the operation of the FBA sensor.
Telemetry is down after a lightning strike caused damage to station uplink.
Station is operating normally.
Telemetry restored after station operator installed new NRTS workstation.
Telemetry down due to workstation hard disk failure.
Station operator replaced PC monitor shipped to him.
New EPROMs installed to correct DSP problem.
Station operator performed procedure to recenter masses of CMG-3T.
The radio telemetry equipment at the station has been replaced. Throughput of all telemetered channels improved immediately.
Telemetry continues to be problematic. New radio equipment was shipped to the station this week.
Station is operating normally but there is a problem with telemetry. A field engineer is on site (2002.04.24).